OpenStudy (anonymous):

answer for this please... Write the equation of the line in slope-intercept form that passes through the points (-2, 6) and (-1, 4). y = -2x + 5 y = 2x - 3 y = 2x + 6 y = -2x + 2

OpenStudy (anonymous):

there are maybe 2 ways to solve this: 1º find the slope of line passing through (-2,6) and (-1,4) and then compare with slopes you got in answers. Select the ones that mach and check if they contain any of the points you need. 2º just construct the equation of the line in slope y-intercept form and look which one it is equal to

OpenStudy (anonymous):

so the answer is y = 2x - 3

OpenStudy (anonymous):

to find the slope of line that pass through points A(a,b) and B(c,d) use this formula: m=d-b/c-a can you do it?

OpenStudy (anonymous):

i dont think so i kinda need help

OpenStudy (anonymous):

ok, in this case A(-2,6) and B(-1,4), then use the formula i gave you befor

OpenStudy (anonymous):

m=4-6/-1-(-2)=-2/1=-2 no?

OpenStudy (anonymous):

so wat could the answer be

OpenStudy (anonymous):

so there are 2 equations with this slope: y = -2x + 5 and y = -2x + 2

OpenStudy (anonymous):

now plug the numbers of one of the points into each equation to see if it matches it

OpenStudy (anonymous):

D is the answer....

OpenStudy (anonymous):

thankssss

OpenStudy (anonymous):

for example: (-1,4) in 1º equation 4 = -2(-1) +5 =8 so doesn't match 2º equation 4=-2(-1)+2 = 4 OK

OpenStudy (anonymous):

oohhh i seeeee nowww so it is D

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@nitz has no patients

OpenStudy (anonymous):

hhahah lol

OpenStudy (anonymous):

what did you say about me@myko?

OpenStudy (anonymous):

@nitz has no patients

OpenStudy (anonymous):

patients??????????????means what?

OpenStudy (anonymous):

to be patient

OpenStudy (anonymous):

i think i spell it right

OpenStudy (anonymous):

patient person

OpenStudy (anonymous):

wait you mean to say that i have no patience ?????????is it????@myko